# DART: An Automated End-to-End Object Detection Pipeline with Data Diversification, Open-Vocabulary Bounding Box Annotation, Pseudo-Label Review, and Model Training

### v1

@@ Abstract
Swift and accurate detection of specified objects is crucial for many industrial applications, such as safety monitoring on construction sites. However, traditional approaches rely heavily on arduous manual annotation and data collection, which struggle to adapt to ever-changing environments and novel target objects. To address these limitations, this paper presents DART, an automated end-to-end pipeline designed to streamline the entire workflow of an object detection application from data collection to model deployment. DART eliminates the need for human labeling and extensive data collection while excelling in diverse scenarios. It employs a subject-driven image generation module (DreamBooth with SDXL) for data Diversification, followed by an Annotation stage where open-vocabulary object detection (Grounding DINO) generates bounding box annotations for both generated and original images. These pseudo-labels are then Reviewed by a large multimodal model (GPT-4o) to guarantee credibility before serving as ground truth to Train real-time object detectors (YOLO). We apply DART to a self-collected dataset of construction machines named Liebherr Product, which contains over 15K high-quality images across 23 categories. The current implementation of DART significantly increases average precision (AP) from 0.064 to 0.832. Furthermore, we adopt a modular design for DART to ensure easy exchangeability and extensibility. This allows for a smooth transition to more advanced algorithms in the future, seamless integration of new object categories without manual labeling, and adaptability to customized environments without extra data collection. The code and dataset are released at https://github.com/chen-xin-94/DART.
@@
journal: Automation in Construction
@@ 1 Introduction
Figure 1: DART, an automated end-to-end pipeline spanning the entire workflow of an object detection application from data collection to model deployment. After preprocessing, DART enters the data Diversification stage, where a subject-driven image generation module (Dreambooth with SDXL) creates diversified images. The Annotation stage (red) utilizes both approved diversified data and original clean data to produce bounding boxes through open-vocabulary object detection (Grounding DINO). These pseudo-labels are then Reviewed by another large multimodal model (GPT-4o) and subsequently serve as ground truth for the Training of real-time object detectors (YOLO).
@@
Despite the above-mentioned remarkable progress in image generation, LMM, OVD, and real-time object detection, each technology still has its prominent strengths and weaknesses. Image generation produces high-quality images but can sometimes not be photorealistic. LMMs excel in semantic understanding of images but struggle with directly handling numerical data. OVD can detect a wide range of object categories but is often too slow. In contrast, real-time object detection is fast but limited to predefined categories. If the strengths of these four technologies can be effectively combined and integrated, it will undoubtedly result in a significantly improved outcome. The proposed DART pipeline (Figure 1) aims to achieve this complementary combination of image generation, LMM, OVD, and real-time object detection, leveraging their strengths and compensating for their weaknesses to create a comprehensive and efficient object detection solution from data collection to model deployment.
@@
Ideally, an object detection application should be capable of detecting any target object both quickly and accurately in diverse environments. However, current technologies do not allow for the simultaneous achievement of open-vocabulary detection, real-time capability, high accuracy, and robustness. The most common approach in the industry is to use real-time object detection to meet the essential requirement for speed while manually collecting and annotating as much data as possible to approach the other three goals [24, 25, 26]. Despite the availability of numerous public datasets, they offer limited assistance in this regard because a majority of specific objects are simply not covered by public datasets. For instance, LVIS [27], one of the most diversified object detection datasets, only includes little more than 1000 entry-level object categories, which is inadequate for many industrial applications. OVD models, on the contrary, are designed to alleviate the object coverage and manual labeling issue. However, they are excessively demanding in terms of computation and memory, making them unsuitable for real-time and on-device applications. On the other hand, the latest image generation tools, such as Stable Diffusion models, are promising alternatives to the inefficient and costly manual collection of more images of the same target object from different scenarios due to their ability for diverse image generation. However, these models lack the ability to mimic the appearance of a specific subject given by visual prompts. Customized frameworks such as DreamBooth could be applied on top for subject-driven image generation instead to address this issue. The generated images can then serve as inputs for OVD to produce bounding boxes, acting as pseudo-labels for training downstream real-time object detection models. However, the results from generative AI can sometimes be overly fanciful, and the quality of bounding boxes produced by OVD can be difficult to guarantee. This necessitates a filtering mechanism to ensure the quality of both generated images and pseudo-labels for the downstream training. This task was previously thought to be exclusive to humans, but now LMMs are well-suited for this task. Although falling short in performing number-related tasks, such as generating bounding box coordinates directly from an input image, they excel in understanding the semantic context of images. As a result, LMM can review the quality of generated images and annotations by solving a designed vision-language task in a c(VQA) manner. To summarize, the foundational advancements in four areas of computer vision collectively form the backbone of the proposed automated end-to-end object detection pipeline DART to fulfill the four objectives of an ideal object detection application. Specifically, DART utilizes subject-driven image generation frameworks for data Diversification to enhance robustness, OVD for ground truth bounding box Annotation to extend object categories without human efforts, LMM for Reviewing generated images and bounding boxes to ensure accuracy and explainability, and Trained real-time object detector as the final output model.
@@
    We propose an automated end-to-end object detection pipeline named DART (Figure 1). It covers the entire workflow of an object detection application, from data collection to model deployment. Specifically, it utilizes subject-driven image generation for data Diversification, open-vocabulary object detection (OVD) for ground truth bounding box Annotation, large multimodal model (LMM) for Reviewing generated images and pseudo-labels, and Trained real-time object detector as the final output.
@@ 2.4 Real-time object detection
Real-time object detection aims to classify and locate objects in real-time with low latency, which is essential for safety-related industrial applications, e.g., on a construction site. This area has witnessed significant advancements in recent years, primarily driven by the iterative development of state-of-the-art YOLO models. Early development of YOLO from v1 to v3 [13, 14, 15] determined the optimal segmentation of the real-time object detection network structure: backbone, neck, and head. YOLOv4 [16] and YOLOv5 [17] changed the backbone to CSPNet [47] and introduced a series of best practices in computer vision regarding training strategy, post-processing, as well as plugin modules. Based on these improvements, YOLOv6 [20] and YOLOv7 [19] are further fine-tuned in label assignment strategy[48] and loss functions [49, 50, 51], and they utilize Repblock [52] and E-ELAN [53] as the core of their backbone respectively. YOLOv8 [21] introduces the C2f building block for effective feature extraction and fusion, along with a comprehensive set of integrated tools, solidifying its position as the go-to choice for real-time object detection. YOLOv9 [22] proposes programmable gradient information (PGI) to refine the training process and a new lightweight network architecture – Generalized Efficient Layer Aggregation Network (GELAN). Finally, the concurrent work YOLOv10 [23] presents an NMS-free training strategy and polishes the mode architecture with holistic efficiency-accuracy-driven design, contributing to its state-of-the-art real-time object detection capabilities. Considering YOLOv8’s ease of use and YOLOv10’s optimal tradeoff between speed and accuracy, this work employs both the YOLOv8 and YOLOv10 families as the real-time object detectors of choice. Although we have selected the best YOLO models, their reliance on predefined object categories restricts their applicability in open scenarios. Therefore, real-time object detection only serves as the last stage in DART, depending on the other three stages to provide high-quality training data.
@@ 3 Methods
The proposed method, DART, is an automated end-to-end object detection pipeline that spans the entire workflow of an object detection application from data collection to model deployment. As illustrated in Figure 1, DART starts with dataset preprocessing, followed by four main stages: data Diversification through subject-driven image generation, bounding box Annotation via OVD, LMM-based Review of generated images and pseudo-labels, as well as real-time object detector Training. The core module (bold text in Figure 1) of each phase is instantiated with state-of-the-art techniques, including the framework DreamBooth with SDXL for subject-driven image generation, the application of Grounding DINO for OVD, the adoption of GPT-4o and InternVL-1.5 for generation quality review, and the training of YOLOv8 and YOLOv10 models for real-time object detection. The following sections provide a detailed breakdown of each component of the proposed DART pipeline.
@@ 3.3 Open-vocabulary bounding box annotation via GroundingDINO
Besides the problem statement and requirement specification, the reminder section for the text prompt is also vital for an optimal response. We find that GPT-4 models tend to be overly strict for bounding box edges and inconsistent when dealing with occluded objects. Consequently, we summarize these common issues in text form (see LABEL:tab:prompt_gpt and other specific requirements for the prompt detailed in Section B.3) and append them at the end of the original text prompt as a reminder. This operation makes GPT-4o much less likely to make the same mistakes.
@@ 4.2 Evaluation metrics
In this work, we adopt average precision (AP) to evaluate the final object detection models. AP first calculates the precision $p$ and recall $r$ at different Intersection over Union (IoU) thresholds $t$ as follows:
@@
where detections are classified as true (T) or false (F) according to their confidence scores and as positive (P) or negative (N) based on the IoU threshold $t$.
Then a modified version of precision-recall curve $\tilde{p}(r)$ is constructed separately for each class $c\in C$:
$\tilde{p}(r,c)=\max_{r^{\prime}\geq r}p(r^{\prime},c),$ (4)
so that the curve will decrease monotonically instead of the zigzag pattern.
Finally, we adopt two commonly used AP metrics: AP50 and AP50-95. The former measures precision at a single IoU threshold of 50%, while the latter averages precision across 10 IoU thresholds ranging from 50% to 95% in increments of 5% ([.50:.05:.95]). Mathematically, AP50-95 can be expressed as follows:
$\text{AP}_{50-95}=\frac{1}{|C|}\sum_{c\in C}\frac{1}{10}\sum_{t\in[.50:.05:.95]}\tilde{p}(r(t),c).$ (5)
AP50-95 is always smaller than AP50 since the precision decreases for a given recall with higher thresholds. However, both metrics exhibit similar trends in diverse and complex datasets, such as LP. This consistent pattern will also be observed in the following tables. Consequently, the following sections will primarily focus on the AP50-95. It is important to note that AP50-95 is also commonly abbreviated as AP, so we will use these two terms interchangeably.
@@ B.3 LMM
GPT-4o receives detailed prompts for bounding box review. The prompts are broken into parts in LABEL:tab:prompt_gpt, with the purposes given on the left and the exact text contents given on the right. First, The system prompt briefly summarizes the bounding box review task and highlights three key metrics: precision, recall, and fit. The user prompt consists of the five remaining sections. The ”task” section outlines the task and objectives in more detail, with {target} referring to the pre-defined class during data collection and {secondary_target} indicating commonly co-occurring classes also recorded during data collection. Note that explicit mention of target types is not mandatory since the boxes in the images already have class labels, and GPT-4o exhibits excellent OCR capability. But it’s still beneficial for corner cases where the label in the image is occluded. Next, the ”questions” section elaborates on the specific questions corresponding to the three main criteria, and the fixed format for output is specified to facilitate subsequent processing in the ”output” segment. In our early experiments, we found that GPT-4 models tend to be overly strict in their judgments regarding the bounding box. To mitigate this, we summarize common issues in the ”relaxation” and ”reminder” sections and attach them to the end of the user prompt. Additionally, our tests with GPT-4v show that including the ”relaxation” and ”reminder” sections as part of the second round of dialogue instead of a single long text is more effective. However, GPT-4o can handle all sections of the user prompt at once.

### v2

@@ Abstract
Accurate real-time object detection is vital across numerous industrial applications, from safety monitoring to quality control. Traditional approaches, however, are hindered by arduous manual annotation and data collection, struggling to adapt to ever-changing environments and novel target objects. To address these limitations, this paper presents DART, an innovative automated end-to-end pipeline that revolutionizes object detection workflows from data collection to model evaluation. It eliminates the need for laborious human labeling and extensive data collection while achieving outstanding accuracy across diverse scenarios. DART encompasses four key stages: (1) Data Diversification using subject-driven image generation (DreamBooth with SDXL), (2) Annotation via open-vocabulary object detection (Grounding DINO) to generate bounding box and class labels (3) Review of generated images and pseudo-labels by large multimodal models (InternVL-1.5 and GPT-4o) to guarantee credibility, (4) Training of real-time object detectors (YOLOv8 and YOLOv10) using the verified data as ground truth. We apply DART to a self-collected dataset of construction machines named Liebherr Product, which contains over 15K high-quality images across 23 categories. The current instantiation of DART significantly increases average precision (AP) from 0.064 to 0.832. Its modular design ensures easy exchangeability and extensibility, allowing for future algorithm upgrades, seamless integration of new object categories, and adaptability to customized environments without manual labeling and additional data collection. The code and dataset are released at https://github.com/chen-xin-94/DART.
@@
journal: Expert Systems with Applications
@@ 1 Introduction
Figure 1: DART, an automated end-to-end pipeline spanning the entire workflow of an object detection application from data collection to model evaluation. After preprocessing, DART enters the data Diversification stage, where a subject-driven image generation module (Dreambooth with SDXL) creates diversified images. The Annotation stage (red) utilizes both approved diversified data and original clean data to produce bounding boxes through open-vocabulary object detection (Grounding DINO). These pseudo-labels are then Reviewed by another large multimodal model (GPT-4o) and subsequently serve as ground truth for the Training of real-time object detectors (YOLO).
@@
Despite the above-mentioned remarkable progress in image generation, LMM, OVD, and real-time object detection, each technology still has its prominent strengths and weaknesses. Image generation produces high-quality images but can sometimes not be photorealistic. LMMs excel in semantic understanding of images but struggle with directly handling numerical data. OVD can detect a wide range of object categories but is often too slow. In contrast, real-time object detection is fast but limited to predefined categories. If the strengths of these four technologies can be effectively combined and integrated, it will undoubtedly result in a significantly improved outcome. The proposed DART pipeline (Figure 1) aims to achieve this complementary combination of image generation, LMM, OVD, and real-time object detection, leveraging their strengths and compensating for their weaknesses to create a comprehensive and efficient object detection solution from data collection to model evaluation.
@@
Ideally, an object detection application should be capable of detecting any target object both quickly and accurately in diverse environments. However, current technologies do not allow for the simultaneous achievement of open-vocabulary detection, real-time capability, high accuracy, and robustness. The most common approach in the industry is to use real-time object detection to meet the essential requirement for speed while manually collecting and annotating as much data as possible to approach the other three goals [24, 25, 26, 27]. Despite the availability of numerous public datasets, they offer limited assistance in this regard because a majority of specific objects are not covered by public datasets. For instance, LVIS [28], one of the most diversified object detection datasets, only includes little more than 1000 entry-level object categories, which is inadequate for many industrial applications. OVD models, on the contrary, are designed to alleviate the object coverage and manual labeling issue. However, they are excessively demanding in terms of computation and memory, making them unsuitable for real-time and on-device applications. On the other hand, the latest image generation tools, such as Stable Diffusion models, are promising alternatives to the inefficient and costly manual collection of more images of the same target object from different scenarios due to their ability for diverse image generation. However, these models lack the ability to mimic the appearance of a specific subject given by visual prompts. Customized frameworks such as DreamBooth could be applied on top for subject-driven image generation instead to address this issue. The generated images can then serve as inputs for OVD to produce bounding boxes, acting as pseudo-labels for training downstream real-time object detection models. However, the results from generative AI can sometimes be overly fanciful, and the quality of bounding boxes produced by OVD can be difficult to guarantee. This necessitates a filtering mechanism to ensure the quality of both generated images and pseudo-labels for the downstream training. This task was previously thought to be exclusive to humans, but now LMMs are well-suited for this task. Although falling short in performing number-related tasks, such as generating bounding box coordinates directly from an input image, they excel in understanding the semantic context of images. As a result, LMM can review the quality of generated images and annotations by solving a designed vision-language task in a Visual Question Answering (VQA) manner. To summarize, the foundational advancements in four areas of computer vision collectively form the backbone of the proposed automated end-to-end object detection pipeline DART to fulfill the four objectives of an ideal object detection application. Specifically, DART utilizes subject-driven image generation frameworks for data Diversification to enhance robustness, OVD for ground truth bounding box Annotation to extend object categories without human efforts, LMM for Reviewing generated images and bounding boxes to ensure accuracy and explainability, and Trained real-time object detector as the final output model.
@@
    We propose an automated end-to-end object detection pipeline named DART (Figure 1). It covers the entire workflow of an object detection application, from data collection to model evaluation. Specifically, it utilizes subject-driven image generation for data Diversification, open-vocabulary object detection (OVD) for ground truth bounding box Annotation, large multimodal model (LMM) for Reviewing generated images and pseudo-labels, and Trained real-time object detector as the final output.
@@ 2.4 Real-time object detection
Real-time object detection aims to accurately classify and locate objects with low latency, which is essential in numerous industrial applications. This area has witnessed significant advancements in recent years, primarily driven by the iterative development of state-of-the-art YOLO models. Early development of YOLO from v1 to v3 [13, 14, 15] determined the optimal segmentation of the real-time object detection network structure: backbone, neck, and head. YOLOv4 [16] and YOLOv5 [17] changed the backbone to CSPNet [48] and introduced a series of best practices in computer vision regarding training strategy, post-processing, as well as plugin modules. Based on these improvements, YOLOv6 [20] and YOLOv7 [19] are further fine-tuned in label assignment strategy[49] and loss functions [50, 51, 52], and they utilize Repblock [53] and E-ELAN [54] as the core of their backbone respectively. YOLOv8 [21] introduces the C2f building block for effective feature extraction and fusion, along with a comprehensive set of integrated tools, solidifying its position as the go-to choice for real-time object detection. YOLOv9 [22] proposes programmable gradient information (PGI) to refine the training process and a new lightweight network architecture – Generalized Efficient Layer Aggregation Network (GELAN). Finally, the concurrent work YOLOv10 [23] presents an NMS-free training strategy and polishes the mode architecture with holistic efficiency-accuracy-driven design, contributing to its state-of-the-art real-time object detection capabilities. Considering YOLOv8’s ease of use and YOLOv10’s optimal tradeoff between speed and accuracy, this work employs both the YOLOv8 and YOLOv10 families as the real-time object detectors of choice. Although we have selected the best YOLO models, their reliance on predefined object categories restricts their applicability in open scenarios. Therefore, real-time object detection only serves as the last stage in DART, depending on the other three stages to provide high-quality training data.
@@ 3 Methods
The proposed method, DART, is an automated end-to-end object detection pipeline that spans the entire workflow of an object detection application from data collection to model evaluation. As illustrated in Figure 1, DART starts with dataset preprocessing, followed by four main stages: data Diversification through subject-driven image generation, bounding box Annotation via OVD, LMM-based Review of generated images and pseudo-labels, as well as real-time object detector Training. The core module (bold text in Figure 1) of each phase is instantiated with state-of-the-art techniques, including the framework DreamBooth with SDXL for subject-driven image generation, the application of Grounding DINO for OVD, the adoption of GPT-4o and InternVL-1.5 for generation quality review, and the training of YOLOv8 and YOLOv10 models for real-time object detection. The following sections provide a detailed breakdown of each component of the proposed DART pipeline.
@@ 3.3 Open-vocabulary bounding box annotation via GroundingDINO
Besides the problem statement and requirement specification, the reminder section for the text prompt is also vital for an optimal response. We find that GPT-4 models tend to be overly strict for bounding box edges and inconsistent when dealing with occluded objects. Consequently, we summarize these common issues in text form (see LABEL:tab:prompt_gpt and other specific requirements for the prompt detailed in Section B.3) and append them at the end of the original text prompt as a reminder. This operation makes GPT-4o much less likely to make the same mistakes. Furthermore, we also utilize Chain-of-Thought (COT) [61] reasoning to enhance the model’s performance. The final text prompt used for pseudo-label review task is detailed in Section B.3.
@@ 4.2 Evaluation metrics
In this work, we utilize the average precision (AP) of the trained object detection models in the last stage as the evaluation metrics for the proposed DART pipeline. AP summarizes the precision-recall (PR) curve into a single value, serving as an approximation of the area under the PR curve. To compute AP, precision $p$ and recall $r$ are first evaluated as follows:
@@
where detections are classified as true (T) or false (F) according to their confidence scores and determined as positive (P) or negative (N) based on the Intersection over Union (IoU) threshold $t$.
Then a modified version of PR curve $\tilde{p}(r,t,c)$ is constructed for each class $c\in C$ given a certain IoU $t$:
$\tilde{p}(r,t,c)=\max_{r^{\prime}\geq r}p(r^{\prime},t,c),$ (4)
so that the curve will decrease monotonically instead of in a zigzag pattern.
[…]
$\text{AP}_{50-95}=\frac{1}{10}\sum_{t\in[.50:.05:.95]}\left(\frac{1}{|C|}\sum_{c\in C}\left(0.01\sum_{r\in[0:.01:1]}\tilde{p}(r,t,c)\right)\right)$ (5)
Note that AP50-95 is no greater than AP50 since the precision decreases for a given recall with higher thresholds. For a well-trained model on diverse and complex datasets, such as LP, both metrics exhibit similar trends. These two consistent patterns will also be observed in the following tables. Consequently, the following sections will primarily focus on the AP50-95. Since AP50-95 is also commonly abbreviated as AP, we will use these two terms interchangeably.
@@ B.3 LMM
GPT-4o receives detailed prompts for bounding box review. The prompts are broken into parts in LABEL:tab:prompt_gpt, with the purposes given on the left and the exact text contents given on the right. First, The system prompt briefly summarizes the bounding box review task and highlights three key metrics: precision, recall, and fit. The user prompt consists of the five remaining sections. The ”task” section outlines the task and objectives in more detail, with {target} referring to the pre-defined class during data collection and {secondary_target} indicating commonly co-occurring classes also recorded during data collection. Note that explicit mention of target types is not mandatory since the boxes in the images already have class labels, and GPT-4o exhibits excellent OCR capability. But it’s still beneficial for corner cases where the label in the image is occluded. Next, the ”questions” section elaborates on the specific questions corresponding to the three main criteria. The output format is specified in the ”output” segment to facilitate subsequent processing. The ”output” section also follows the guidelines of COT to enhance the model’s performance. In our early experiments, we found that GPT-4 models tend to be overly strict in their judgments regarding the bounding box. To mitigate this, we summarize common issues in the ”relaxation” and ”reminder” sections and attach them to the end of the user prompt. Additionally, our tests with GPT-4v show that including the ”relaxation” and ”reminder” sections as part of the second round of dialogue instead of a single long text is more effective. However, GPT-4o can handle all sections of the user prompt at once.